Ultraviolet light of wavelength Lambda 1 and Lambda 2 when allowed to fall on a hydrogen atoms in their Ground state is found to liberate electrons with kinetic energy 1.8ev and 4ev respectively.find the value of lambda 1/lambda 2 (L1/L2)

Grade:12

2 Answers

Amarjit
6 years ago
Since electrons are liberated hence the total energoes must be 15.4eV and 17.6eV respectively.Now using the formula E=hc/lambda we will get lambda in both the cases keeping hc= 12400( lamba in the order of angstrom order).Now just take their ratio.
Ankit Khandelwal
2 years ago
Since electron travel from ground state to infinity so the energy liberated is 13.6 ev + kinetic energy
                                                                                                                                         13.6   + 1.8 =15.4ev
                                                                                                                                          13.6  +  4 =  17.6 ev
Now we know E=hc/lamda(in angstrom)
we want lambda1/lambda2= 17.6/15.4 = 8/7